OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (-2, -3) and is parallel to the line y = x + 9?

OpenStudy (tpaulus):

y+3=x+2 y=x-1 y-x=-1

OpenStudy (tpaulus):

Slope = 1 Use point slope form then convert.

OpenStudy (anonymous):

Since the line is parallel to the line y=x+9 then they have the same slope 1 ! SO the line has an equation like this : \[y=x+b\] Now, to find b we have (-2,-2) is from the line , so : \[-2=-2+b\] and then : \[b=0\] So the equation is : \[y=x\]

OpenStudy (anonymous):

Thank you!

OpenStudy (tpaulus):

@Noura11 Your answer doesn't go through the point(-2,-3)

OpenStudy (anonymous):

Ah I I have done a mistake ! I should say : Now, to find b we have (-2,-3) is from the line , so : −3=−2+b and then : b=-3+2=-1 So the equation is : y=x-1

OpenStudy (anonymous):

Ok thanks!

OpenStudy (tpaulus):

No problem, simple algebra